OpenStudy (anonymous):

find all real zeros for the polynomial function. x^3-2x^2- 9x+18

OpenStudy (anonymous):

All of the factors of the constant divided by the coefficient of the highest degree. In this case +-18/+-1 = +18,-18,-9,+9,+2,-2, +3, -3, and +6, -6

OpenStudy (anonymous):

What is the constant?

OpenStudy (anonymous):

the last number in the polynomial

OpenStudy (anonymous):

So it's just 18/1

Directrix (directrix):

@katieeejay This polynomial will factor. You can use the Rational Root Theorem but it is not necessary. x^3-2x^2- 9x+18=0 x^2( x - 2) - 9 ( x - 2) = 0 (x - 2) * (x^2 - 9) = 0 (x - 2) * (x + 3)* (x - 3) = 0 Zero Product Property Sove for x to get the roots.